Nick Brewer: Difference between revisions

Jump to navigation Jump to search
no edit summary
No edit summary
No edit summary
'''<big>5/23/14</big>'''
 
Today the cavity is working great. There were a few small peaks between the 0,0 peaks but they were small and I was able to get rid of them by adjusting the mirrors and mode matching lenses. Yesterday's problems might have been a laser issue due to the weather. This morning the humidity was 30%, I'm not sure what it was yesterday but I should keep an eye on how that affects how well the cavity is behaving.
 

Navigation menu